solve the system by substitution 5x+3y=-17, y=3x-15
Answer by checkley75(3666)

Y=3X-15 NOW WE SUBSTITUTE (3X-15) FOR Y IN THE OTHERE EQUATION & SOLVE FOR X
5X+3(3X-15)=-17
5X+9X-45=-17
14X=-17+45
14X=28
X=28/14
X=2 ANSWER. NOW SUBSTITUTE 2 FOR X IN THE FIRST EQUATION & SOLVE FOR Y
Y=3*2-15
Y=6-15
Y=-9 ANSWER.
PROOF
5*2+3*-9=-17
10-27=-17
-17=-17